OpenStudy (anonymous):

Two Martians, Splott and Fizzle, have solved the equation 2x + 4 = –3x + 14. Examine the work of Splott and Fizzle. Identify any errors in the Martians’ calculations and explain, using complete sentences, what corrections they should make.

OpenStudy (texaschic101):

need to see what they did

OpenStudy (carolina4567):

I had to do this as well what module is it again ?

OpenStudy (carolina4567):

Splott is wrong in 5x + 4 + 4 = 14 + 4 he should have subtracted 4.

OpenStudy (anonymous):

module 1

OpenStudy (cwrw238):

funny - I was born in Splott!!

OpenStudy (carolina4567):

New answer for splott is = 10/5 New answer for fizzle is = -2 = x

OpenStudy (carolina4567):

Using complete sentences, prove to Splott and Fizzle that your function is a legitimate function. In mathematics, a function is a relation between a set of inputs and a set of permissible outputs with the property that each input is related to exactly one output. Which would mean that for any x value there is a y value.

OpenStudy (carolina4567):

3rd question is f(x) = 10x + 6 First we set up f(3) = 10*3 +6 Then we mutiply 10 x 3 = 30+6 we get 30 then we add them up together = 36 and we get our final answer

OpenStudy (anonymous):

Create your own function to teach the Martians about functions. Your function must contain at least two different operations. i dont know how to do functions

OpenStudy (anonymous):

jking i got that

OpenStudy (carolina4567):

f(x) = 10x+6 Basically the inverse is to find x first so....... we do the opposite of what was done before f(x) - 6 = 10x f(x)-6 ------ = x 10 f(x)= x-6 It would look like this ____ 10

OpenStudy (carolina4567):

And last h(x)= 2x - 1 h(x) = 2(10x+6) - 1 h(x) = 20x + 12 - 1 h(g(x)) = 20x + 11 f(x) = 10x + 6 f(x) = 10(2x - 1) +6 f(x) = 20x -10 + 6 f(x) = 20x - 4 x = 2

OpenStudy (carolina4567):

Ur welcome :)

OpenStudy (anonymous):

Using your function, explain to the Martians how to solve for f(3). Show your work and explain each step using complete sentences. Using complete sentences, describe to the Martians how to find the inverse of your function. The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2

OpenStudy (carolina4567):

All ur answers are above ^^^^^^^^^^^